# A Survey of Text Classification Under Class Distribution Shift

### v2

@@ Abstract
The basic underlying assumption of machine learning (ML) models is that the training and test data are sampled from the same distribution. However, in daily practice, this assumption is often broken, i.e. the distribution of the test data changes over time, which hinders the application of conventional ML models. One domain where the distribution shift naturally occurs is text classification, since people always find new topics to discuss. To this end, we survey research articles studying open-set text classification and related tasks. We divide the methods in this area based on the constraints that define the kind of distribution shift and the corresponding problem formulation, i.e. learning with the Universum, zero-shot learning, and open-set learning. We next discuss the predominant mitigation approaches for each problem setup. Finally, we identify several future work directions, aiming to push the boundaries beyond the state of the art. Interestingly, we find that continual learning can solve many of the issues caused by the shifting class distribution. We maintain a list of relevant papers online [^1] https://github.com/Eduard6421/Open-Set-Survey.
A Survey of Text Classification Under Class Distribution Shift
Adriana Valentina Costache, Silviu Florin Gheorghe, Eduard Gabriel Poesina, Paul Irofti  and Radu Tudor Ionescuthanks: Corresponding author: raducu.ionescu@gmail.com. Equal contribution. University of Bucharest Bucharest, Romania
@@ 1 Introduction
The primary assumption of any machine learning (ML) model is that the data is independent and identically distributed (IID) Vapnik (1995). In learning theory, the IID assumption plays a critical role, as it guarantees the generalization capacity of models, provided that sufficient training data is available, and that the hypothesis class is not too large. However, this assumption does not always hold in daily practice. One such example is text categorization by topic, where new topics naturally emerge as the interest of people changes over time. For example, journalists started publishing news articles about the COVID-19 pandemic only after its outbreak began in December 2019222https://en.wikipedia.org/wiki/COVID-19_pandemic. Hence, if a classifier is trained in a closed-world setup (a scenario where it is presumed that every instance encountered by the model belongs to a class that is present in the training data), there is a high risk that the respective model will encounter instances from classes that were not present in the training data, during operation. Such instances will be misclassified as belonging to one of the training classes, degrading the overall performance of the ML system, and implicitly reducing user satisfaction.
@@
A large body of works, e.g. Geng et al. (2021); Salehi et al. (2022); Vaze et al. (2021); Yang et al. (2024), studies the adaptation and application of ML models when the class distribution changes from training time to inference time. One of the first articles to highlight this problem is that of Scheirer et al. (2013). The authors define the open-set risk as the sum of the empirical risk and the open-space risk, the latter being defined as the risk of misclassifying instances of unknown classes as belonging to a known class. Over time, researchers attempted to minimize the open-set risk by addressing a broad range of formulations of the class distribution shift problem. These formulations can be mainly structured according to the class categories proposed by Geng et al. (2021).
@@
KUCs Classes with available training examples, but for which there are no class labels. Classification with background / Universum class Dhamija et al. (2018); Hendrycks et al. (2019); Lee et al. (2018); Liu et al. (2020)
UKCs Classes that are known to exist, but for which we have no examples during training. Zero-shot learning Gera et al. (2022); Meng et al. (2020, 2022); Sanh et al. (2022); Yin et al. (2019); Zhang et al. (2024)
UUCs Classes whose existence is unknown during training and for which there is no training data. Open-set learning (and discovery) Chen et al. (2024, 2023); Kim et al. (2022); Walkowiak et al. (2019a, 2020)
@@
We present the class categories and the corresponding problem formulations in Figure 1. We further define the class categories and provide representative studies for each problem formulation in Table 1. KKCs lead to the conventional supervised learning paradigm, which is well-studied and results often surpass human-level performance Cozma et al. (2018); Tedeschi et al. (2023). Frameworks that deal with KUCs typically use a background (a.k.a. Universum) class Dhamija et al. (2018); Weston et al. (2006), essentially extending the supervised learning setup with a new category where all the examples that do not belong to any of the KKCs are placed. The zero-shot learning paradigm Chaudhary et al. (2024); Pourpanah et al. (2023); Yin et al. (2019) aims to deal with UKCs, namely with classes that are known in advance (at training time), but for which there are no data samples. Zero-shot learning aims to classify data samples into UKCs, but the main limitation is that the set of UKCs is fixed. As a consequence, dealing with UUCs is not possible in the zero-shot learning setup. Open-set learning Geng et al. (2021); Scheirer et al. (2013) aims to partially address this challenge by identifying UUCs without having prior information about such classes during training. Sometimes, UUCs are identified via a two-stage pipeline that combines an outlier detection (a.k.a. novelty detection) model Chen et al. (2023); Walkowiak et al. (2018, 2020) and a supervised model trained on KKCs. However, unlike zero-shot learning methods, open-set methods do not aim to classify the data samples into UUCs. A more comprehensive setup is proposed in Zheng et al. (2022), where the authors aim not only to detect samples belonging to UUCs, but also to classify them. This framework, called open-set learning and discovery, can be seen as a generalization of zero-shot learning, where the set of UKCs can be updated during inference, therefore transforming it into a set of UUCs.
@@
To date, open-set classification and the related tasks discussed above were mostly studied in the vision domain, where various tasks have been explored, such as object recognition Kong and Ramanan (2021); Scheirer et al. (2013); Vaze et al. (2021), semantic segmentation Cen et al. (2021); Oliveira et al. (2021), object detection Dhamija et al. (2020); Liu et al. (2024); Zheng et al. (2022), and video anomaly detection Acsintoae et al. (2022); Wu et al. (2024), among others. Comparatively less attention has been dedicated to this family of tasks in the text domain Chen et al. (2024); Fei and Liu (2016); Kim et al. (2022). However, class distribution shift is a prevalent phenomenon in text classification. Aside from the example provided earlier about the changing topics over time, there are many other natural language processing (NLP) tasks where the class distribution can shift. For example, in authorship identification, new authors can emerge overtime, so an authorship identification model needs to update the list of potential authors. Another task affected by class distribution shift is intent detection in conversational AI. For instance, a chatbot trained to recognize intents such as “book flight” or “check weather” may encounter new intents such as “play music” at test time. Similarly, in named entity recognition, a system trained to recognize entities such as “person” and “location” might face new entity types such as “event” or “disease” during operation. Despite the clear likelihood of encountering new classes during inference in various NLP tasks, to the best of our knowledge, there is no survey on open-set text classification. To this end, we review articles that address the class distribution shift in the text domain. We further propose future work directions to address the challenges of open-set learning and discovery in NLP tasks, most of which stem from the continual learning paradigm Wang et al. (2024).
@@ 2 Notations
We use the following notations to define the various task formulations covered in our survey. Let $x\in\mathcal{X}$ represent a text sample from the data space $\mathcal{X}$, and $y\in\mathcal{Y}$ a label from the label space $\mathcal{Y}$. The label space is divided into four disjoint subsets denoted as known known classes ($C^{kk}$), known unknown classes ($C^{ku}$) unknown known classes ($C^{uk}$), and unknown unknown classes ($C^{uu}$), such that $C^{kk}\cup C^{ku}\cup C^{uk}\cup C^{kk}=\mathcal{Y}$ and the intersection between any two subsets is the empty set, e.g. $C^{kk}\cap C^{uk}=\emptyset$.
@@ 3 Learning with Background Class
Formal definition. Learning with background / Universum class aims to detect samples from $C^{ku}$, while having samples from both $C^{kk}$ and $C^{ku}$. Formally, the training data is defined as $\mathcal{D}=\{(x,y)\mid x\in\mathcal{X},y\in C^{kk}\cup C^{ku}\}$, while the test data is defined as $\mathcal{T}=\{(x,y)\mid x\in\mathcal{X},y\in C^{kk}\cup C^{ku}\}$. In this setup, a classifier is defined as $h:\mathcal{X}\to C^{kk}\cup\{\text{background}\}$, i.e. for any sample that belongs to $C^{ku}$, the model should label the respective sample as background.
@@
Trends and approaches. The problem of learning with a background class has been addressed through proxy tasks such as out-of-distribution (OOD) detection or outlier rejection Dhamija et al. (2018). These approaches improve the reliability of the model by managing uncertainty and rejecting inputs outside the distribution of KKCs. Research in this area has been predominantly driven by computer vision studies, whereas text classification remains comparatively underexplored. However, many techniques can be easily adapted to NLP tasks by replacing image-based feature extractors with robust language encoders.
@@
A straightforward approach to detect background samples involves training a classifier to refine the decision boundary between the original dataset and a curated background dataset. This technique is commonly employed in computer vision tasks Mullapudi et al. (2021), leveraging the abundance of background images available, e.g. iNaturalist-BG, Places-BG, etc.
@@
Another common approach is to employ custom regularization techniques for background-class learning. For Support Vector Machines (SVMs), Weston et al. (2006) introduce a penalty term that incorporates the distance between the decision boundary and background-class (Universum-class) data points. Their method is based on the principle that Universum-class samples should lie near the decision boundary, reflecting real-world uncertainty and ambiguity in classification. This regularization technique enforces a decision-making process that accounts for inherent uncertainty in the data. Dhamija et al. (2018) introduce a more generic regularization based on two loss functions. The Entropic Open-Set Loss aims to reduce entropy for KKCs, while maximizing entropy for KUCs, ensuring that the model treats unfamiliar data with the highest possible uncertainty. The Objectosphere Loss strengthens the separation between KKCs and KUCs by increasing the feature magnitudes for KKC samples, while reducing them for KUCs. These loss functions create a more distinct feature space, improving the ability of the model to reject out-of-distribution inputs.
@@
Developing novel loss functions represents a promising line of work, considered by other researchers as well. For instance, Liu et al. (2020) propose energy-based models as a viable solution for background sample detection, where the main contribution is a loss function designed to include both a standard classification loss and a model-specific penalty term. This additional term penalizes high-energy representations of KKC samples, while suppressing low-energy assignments for KUC samples, ensuring a clear energy-based separation between the two categories. Hendrycks et al. (2019) introduce the concept of outlier exposure, which involves training models on an auxiliary dataset containing OOD samples. Their objective function combines the original classification loss with an additional penalty that makes the model assign low confidence to out-of-distribution inputs. For softmax-based classification tasks, this approach encourages predictions for outliers to follow a uniform distribution. In density estimation tasks, the outlier exposure loss often incorporates a margin ranking objective, ensuring that in-distribution samples consistently receive higher log probabilities than outliers. Lee et al. (2018) propose a method to train neural networks for better OOD detection, while maintaining their classification accuracy. The contributions refer to introducing a confidence loss, leveraging a generative adversarial network (GAN) Goodfellow et al. (2014) for OOD sample generation, and combining these into a joint training framework. The confidence loss combines the standard cross-entropy loss with a KL-divergence term penalizing confident predictions for OOD samples, which are compared with a uniform distribution. The OOD samples are generated by a GAN, such that the samples lie close to the boundary of the in-distribution data. The authors train the classifier and the GAN jointly.

### v3

# Text Classification Under Class Distribution Shift: A Survey

@@ Abstract
The basic underlying assumption of machine learning (ML) models is that the training and test data are sampled from the same distribution. However, in daily practice, this assumption is often broken, i.e. the distribution of the test data changes over time, which hinders the application of conventional ML models. One domain where the distribution shift naturally occurs is text classification, since people always find new topics to discuss. To this end, we survey research articles studying open-set text classification and related tasks. We divide the methods in this area based on the constraints that define the kind of distribution shift and the corresponding problem formulation, i.e. learning with the Universum, zero-shot learning, and open-set learning. We next discuss the predominant mitigation approaches for each problem setup. We further identify several future work directions, aiming to push the boundaries beyond the state of the art. Finally, we explain how continual learning can solve many of the issues caused by the shifting class distribution. We maintain a list of relevant papers at https://github.com/Eduard6421/Open-Set-Survey.
Text Classification Under Class Distribution Shift: A Survey
Adriana-Valentina Costache1,∗, Silviu-Florin Gheorghe1,∗, Eduard Poesina1,∗, Paul Irofti1, Radu Tudor Ionescu1,⋄ 1Department of Computer Science, University of Bucharest, Romania Equal contribution. raducu.ionescu@gmail.com
@@ 1 Introduction
The primary assumption of any machine learning (ML) model is that the data is independent and identically distributed (IID) (Vapnik, 1995). In learning theory, the IID assumption plays a critical role, as it guarantees the generalization capacity of models, provided that sufficient training data is available, and that the hypothesis class is not too large. However, this assumption does not always hold in daily practice. One such example is text categorization by topic, where new topics naturally emerge as the interest of people changes over time. For example, journalists started publishing news articles about the COVID-19 pandemic only after its outbreak began in December 2019111https://en.wikipedia.org/wiki/COVID-19_pandemic. Hence, if a classifier is trained in a closed-world setup (a scenario where it is presumed that every instance encountered by the model belongs to a class that is present in the training data), there is a high risk that the respective model will encounter instances from classes that were not present in the training data, during operation. Such instances will be misclassified as belonging to one of the training classes, degrading the overall performance of the ML system, and implicitly reducing user satisfaction.
@@
KUCs Classes with available training examples, but for which there are no class labels. Classification with background / Universum class (Dhamija et al., 2018; Hendrycks et al., 2019; Lee et al., 2018; Liu et al., 2020)
UKCs Classes that are known to exist, but for which we have no examples during training. Zero-shot learning (Yin et al., 2019; Gera et al., 2022; Meng et al., 2020, 2022; Sanh et al., 2022; Zhang et al., 2024)
UUCs Classes whose existence is unknown during training and for which there is no training data. Open-set learning (and discovery) (Chen et al., 2023; Walkowiak et al., 2020; Chen et al., 2024; Kim et al., 2022; Walkowiak et al., 2019a)
@@
A large body of works, e.g. (Geng et al., 2021; Salehi et al., 2022; Vaze et al., 2021; Yang et al., 2024), studies the adaptation and application of ML models when the class distribution changes from training time to inference time. One of the first articles to highlight this problem is that of Scheirer et al. (2013). The authors define the open-set risk as the sum of the empirical risk and the open-space risk, the latter being defined as the risk of misclassifying instances of unknown classes as belonging to a known class. Over time, researchers attempted to minimize the open-set risk by addressing a broad range of formulations of the class distribution shift problem. These formulations can be mainly structured according to the class categories proposed by Geng et al. (2021).
@@
We present the class categories and the corresponding problem formulations in Figure 1. We further define the class categories and provide representative studies for each problem formulation in Table 1. KKCs lead to the conventional supervised learning paradigm, which is well-studied and results often surpass human-level performance (Cozma et al., 2018; Tedeschi et al., 2023). Frameworks that deal with KUCs typically use a background (a.k.a. Universum) class (Dhamija et al., 2018; Weston et al., 2006), essentially extending the supervised learning setup with a new category where all the examples that do not belong to any of the KKCs are placed. The zero-shot learning paradigm (Yin et al., 2019; Chaudhary et al., 2024; Pourpanah et al., 2023) aims to deal with UKCs, namely with classes that are known in advance (at training time), but for which there are no data samples. Zero-shot learning aims to classify data samples into UKCs, but the main limitation is that the set of UKCs is fixed. As a consequence, dealing with UUCs is not possible in the zero-shot learning setup. Open-set learning (Geng et al., 2021; Scheirer et al., 2013) aims to partially address this challenge by identifying UUCs without having prior information about such classes during training. Sometimes, UUCs are identified via a two-stage pipeline that combines an outlier detection (a.k.a. novelty detection) model (Chen et al., 2023; Walkowiak et al., 2020, 2018) and a supervised model trained on KKCs. However, unlike zero-shot learning methods, open-set methods do not aim to classify the data samples into UUCs. A more comprehensive setup is proposed in (Zheng et al., 2022), where the authors aim not only to detect samples belonging to UUCs, but also to classify them. This framework, called open-set learning and discovery, can be seen as a generalization of zero-shot learning, where the set of UKCs can be updated during inference, therefore transforming it into a set of UUCs.
@@
To date, open-set classification and the related tasks discussed above were mostly studied in the vision domain, where various tasks have been explored, such as object recognition (Vaze et al., 2021; Scheirer et al., 2013; Kong and Ramanan, 2021), semantic segmentation (Cen et al., 2021; Oliveira et al., 2021), object detection (Zheng et al., 2022; Dhamija et al., 2020; Liu et al., 2024), and video anomaly detection (Acsintoae et al., 2022; Wu et al., 2024), among others. Comparatively less attention has been dedicated to this family of tasks in the text domain (Chen et al., 2024; Kim et al., 2022; Fei and Liu, 2016). However, class distribution shift is a prevalent phenomenon in text classification. Aside from the example provided earlier about the changing topics over time, there are many other natural language processing (NLP) tasks where the class distribution can shift. For example, in authorship identification, new authors can emerge over time, so an authorship identification model needs to update the list of potential authors. Another task affected by class distribution shift is intent detection in conversational AI. For instance, a chat bot trained to recognize intents such as “book flight” or “check weather” may encounter new intents such as “play music” at test time. Similarly, in named entity recognition, a system trained to recognize entities such as “person” and “location” might face new entity types such as “event” or “disease” during operation. Despite the clear likelihood of encountering new classes during inference in various NLP tasks, to the best of our knowledge, there is no survey on open-set text classification. To this end, we review articles that address the class distribution shift in the text domain. We further propose future work directions to address the challenges of open-set learning and discovery in NLP tasks, most of which stem from the continual learning paradigm (Wang et al., 2024).
It is perhaps important to note that class distribution shift is a particular kind of dataset shift. The dataset shift problem was studied by Moreno-Torres et al. (2012), who meticulously categorize the various types of dataset shift. However, the authors did not specifically include class distribution shift as an interesting scenario, but they recognize the existence of dataset shift scenarios that “are so hard that we currently consider them impossible to solve” (Moreno-Torres et al., 2012). Since 2012, the state of research in machine learning has changed significantly, and many problems that were thought as being impossible to solve are now actively studied. For instance, there are several recent surveys, e.g. (Geng et al., 2021; Parmar et al., 2023; Zhu et al., 2024), that focus on the open-world learning problem, where new classes can emerge during testing. Yet, most of the studies covered by these surveys come from computer vision and image processing. To the best of our knowledge, we are the first to focus specifically on class distribution shift in natural language processing.
@@ 2 Notations and Definitions
Notations. We use the following notations to define the various task formulations covered in our survey. Let $x\in\mathcal{X}$ represent a text sample from the data space $\mathcal{X}$, and $y\in\mathcal{Y}$ a label from the label space $\mathcal{Y}$. The label space is divided into four disjoint subsets denoted as known known classes ($C^{kk}$), known unknown classes ($C^{ku}$), unknown known classes ($C^{uk}$), and unknown unknown classes ($C^{uu}$), such that $C^{kk}\cup C^{ku}\cup C^{uk}\cup C^{kk}=\mathcal{Y}$ and the intersection between any two subsets is the empty set, e.g. $C^{kk}\cap C^{uk}=\emptyset$.
Learning with background class. Learning with background / Universum class aims to detect samples from $C^{ku}$, while having samples from both $C^{kk}$ and $C^{ku}$. Formally, the training data is defined as $\mathcal{D}=\{(x,y)\mid x\in\mathcal{X},y\in C^{kk}\cup C^{ku}\}$, while the test data is defined as $\mathcal{T}=\{(x,y)\mid x\in\mathcal{X},y\in C^{kk}\cup C^{ku}\}$. In this setup, a classifier is defined as $h:\mathcal{X}\to C^{kk}\cup\{\text{background}\}$, i.e. for any sample that belongs to $C^{ku}$, the model should label the respective sample as background.
Zero-shot text classification. Zero-shot learning aims to recognize samples from $C^{uk}$ using knowledge transferred from $C^{kk}$. Formally, the training data is defined as $\mathcal{D}=\{(x,y)\mid x\in\mathcal{X},y\in C^{kk}\}$, while the test data is defined as set $\mathcal{T}=\{(x,y)\mid x\in\mathcal{X},y\in C^{uk}\}$. Knowledge is usually transferred via an auxiliary semantic space $A$, e.g. word embeddings or attribute vectors, such that $f:\mathcal{X}\to A$ and $g:A\to\mathcal{Y}$. The mapping $f$ projects features into the semantic space, while $g$ maps semantic representations to class labels.
Open-set text classification and discovery. Open-set learning aims to classify samples from $C^{kk}$, while rejecting samples from $C^{uu}$ encountered during inference. The training set is defined as $\mathcal{D}=\{(x,y)\mid x\in\mathcal{X},y\in C^{kk}\}$, while the test set is given by $\mathcal{T}=\{(x,y)\mid x\in\mathcal{X},y\in C^{kk}\cup C^{uu}\}$. An open-set learning model is defined as $h:\mathcal{X}\to C^{kk}\cup\{\text{unknown}\}$, i.e. for any sample that belongs to $C^{uu}$, the model should label the respective sample as unknown. In the open-set learning and discovery setup, the model is defined as $h:\mathcal{X}\to C^{kk}\cup C^{uu}$, i.e. the model must be able to classify any data sample, regardless of the fact that the sample belongs to $C^{kk}$ or $C^{uu}$. To classify samples into $C^{uu}$, classes belonging to $C^{uu}$ must be discovered during inference.
3 Optimization Objectives
Starting from a model whose parameters $\theta$ are obtained by training on a given dataset $\mathcal{X}$ with labels $\mathcal{Y}$, our general optimization problem is:
$\begin{split}\theta^{\star}&=\arg\min_{\theta}\mathcal{L}\left(\theta,\mathcal{X},\mathcal{Y}\right)\\ &=\arg\min_{\theta}\mathbb{E}_{(x,y)\sim(\mathcal{X},\mathcal{Y})}\left[f_{\theta}(x,y)\right].\end{split}$ (1)
We recover below the particular learning frameworks by giving specific loss formulations w.r.t. the class types defined in Table 1. The supervised cross-entropy on KKCs is:
$\mathcal{L}_{\text{KKC}}(\theta,\mathcal{X},C^{kk})\!\!=\!\!\mathbb{E}_{(x,y)\sim(\mathcal{X},C^{kk})}\!\left[-\!\log p_{\theta}(y\!\mid\!x)\right]\!.$ (2)
The binary cross-entropy on KKCs vs. KUCs is:
$\begin{split}\!\!\!\mathcal{L}_{\text{KUC}}&(\theta,\mathcal{X},C^{kk}\cup C^{ku})\!\!=\!\!\mathbb{E}_{(x,y)\sim(\mathcal{X},C^{kk}\cup C^{ku})}\\ &\left[-y\!\cdot\!\log f_{\theta}(x)\!+\!(1\!-\!y)\!\cdot\!\log(1\!-\!f_{\theta}(x))\right]\!.\end{split}$ (3)
The contrastive learning objective on UKCs is:
$\begin{split}\mathcal{L}_{\text{UKC}}&(\theta,\mathcal{X},C^{uk})\!=\!\mathbb{E}_{(x,y)\sim\mathcal{(}\mathcal{X},C^{uk})}\\ &\left[-\!\log\frac{\exp\left(\text{sim}(f_{\theta}(x),s_{y})\right)}{\sum_{c\in\mathcal{C}^{uk}}\exp\left(\text{sim}(f_{\theta}(x),s_{c})\right)}\right]\!,\end{split}$ (4)
where $s_{y}$ is a prototype vector (embedding) for class $y$, which can be obtained by passing the class names through the model $f_{\theta}$, and sim is a similarity measure (e.g. the cosine similarity). Finally, the supervised cross-entropy on UUCs is:
$\mathcal{L}_{\text{UUC}}(\theta,\mathcal{X},C^{uu})\!\!=\!\!\mathbb{E}_{(x,y)\sim(\mathcal{X},C^{uu})}\!\left[-\!\log p_{\theta}(y\!\mid\!x)\right]\!,$ (5)
and the label set is computed as $C^{uu}\!=\!\{y\mid y\!=\!g_{\omega}(x^{\prime}),\forall x^{\prime}\in\mathcal{I}(h_{\phi}(x))\}$, where $h_{\phi}$ is an outlier detection method (e.g. Isolation Forrest) and $\mathcal{I}$ is an unsupervised outlier indicator function that selects outlier samples $x^{\prime}$, which are further given as input to a clustering method $g_{\omega}$ (e.g. k-means).
Further, we can integrate all learning frameworks into a joint optimization objective:
$\mathcal{L}_{\text{joint}}\!=\!\mathcal{L}_{\text{KKC}}\!+\!\alpha\!\cdot\!\mathcal{L}_{\text{KUC}}\!+\!\beta\!\cdot\!\mathcal{L}_{\text{UKC}}\!+\!\gamma\!\cdot\!\mathcal{L}_{\text{UUC}},$ (6)
where $\alpha$, $\beta$ and $\gamma$ are scalar values that can be used to control the optimization within a particular learning framework, as follows:
  • Learning with background class: $\alpha\!>\!0$, $\beta\!=\!0$, $\gamma\!=\!0$.
  • Zero-shot learning: $\alpha\!=\!0$, $\beta\!>\!0$, $\gamma\!=\!0$.
  • Open-set learning and discovery: $\alpha\!=\!0$, $\beta\!=\!0$, $\gamma\!>\!0$.
@@ 4 Learning with Background Class
The problem of learning with a background class has been addressed through proxy tasks such as out-of-distribution (OOD) detection or outlier rejection (Dhamija et al., 2018). These approaches improve the reliability of the model by managing uncertainty and rejecting inputs outside the distribution of KKCs. Research in this area has been predominantly driven by computer vision studies, whereas text classification remains comparatively underexplored. However, many techniques can be easily adapted to NLP tasks by replacing image-based feature extractors with robust language encoders.
@@
A straightforward approach to detect background samples involves training a classifier to refine the decision boundary between the original dataset and a curated background dataset. This technique is commonly employed in computer vision tasks (Mullapudi et al., 2021), leveraging the abundance of background images available, e.g. iNaturalist-BG, Places-BG, etc.
@@
Another common approach is to employ custom regularization techniques for background-class learning. For Support Vector Machines (SVMs), Weston et al. (2006) introduce a penalty term that incorporates the distance between the decision boundary and background-class (Universum-class) data points. Their method is based on the principle that Universum-class samples should lie near the decision boundary, reflecting real-world uncertainty and ambiguity in classification. This regularization technique enforces a decision-making process that accounts for inherent uncertainty in the data. Expanding this idea, Zhou et al. (2023) propose a learning method that distinctively treats the classes of interest and the Universum class. They introduce closed decision boundaries for the classes of interest and define the space outside these boundaries as belonging to the Universum. To determine whether a sample belongs to a class of interest or to the Universum class, the authors propose a probability estimation based on inter-class rules. This approach assesses whether a sample lies predominantly within the closed decision boundary associated with a class of interest. If the sample does not clearly fit into one of the classes, it is considered to belong to the Universum. Dhamija et al. (2018) introduce a more generic regularization based on two loss functions. The Entropic Open-Set Loss aims to reduce entropy for KKCs, while maximizing entropy for KUCs, ensuring that the model treats unfamiliar data with the highest possible uncertainty. The Objectosphere Loss strengthens the separation between KKCs and KUCs by increasing the feature magnitudes for KKC samples, while reducing them for KUCs. These loss functions create a more representative feature space, improving the ability of the model to reject out-of-distribution inputs.
@@
Developing novel loss functions represents a promising line of work, considered by other researchers as well. For instance, Liu et al. (2020) propose energy-based models as a viable solution for background sample detection, where the main contribution is a loss function designed to include both a standard classification loss and a model-specific penalty term. This additional term penalizes high-energy representations of KKC samples, while suppressing low-energy assignments for KUC samples, ensuring a clear energy-based separation between the two categories. Hendrycks et al. (2019) introduce the concept of outlier exposure, which involves training models on an auxiliary dataset containing OOD samples. Their objective function combines the original classification loss with an additional penalty that makes the model assign low confidence to out-of-distribution inputs. For softmax-based classification tasks, this approach encourages predictions for outliers to follow a uniform distribution. In density estimation tasks, the outlier exposure loss often incorporates a margin ranking objective, ensuring that in-distribution samples consistently receive higher log probabilities than outliers. Lee et al. (2018) propose a method to train neural networks for better OOD detection, while maintaining their classification accuracy. The contributions refer to introducing a confidence loss, leveraging a generative adversarial network (GAN) (Goodfellow et al., 2014) for OOD sample generation, and combining these into a joint training framework. The confidence loss combines the standard cross-entropy loss with a KL-divergence term penalizing confident predictions for OOD samples, which are compared with a uniform distribution. The OOD samples are generated by a GAN, such that the samples lie close to the boundary of the in-distribution data. The authors train the classifier and the GAN jointly. A complementary direction is explored by Hu and Khan (2021), who develop a loss function specifically designed to calibrate uncertainty in both in-distribution and OOD settings. Their framework employs evidential neural networks and adds two regularization terms: one penalizes high uncertainty for in-distribution samples, and the other rewards maintaining high uncertainty OOD samples. By explicitly modeling and balancing predictive confidence, their approach improves classification reliability under class distribution shift.